### v3

@@ 5.2 No Approach Dominates the Other
A formal proof can be found in online Appendix B.3. It proceeds by constructing examples where child and family utilites are mismatched, causing the utility gap between child-optimal and family-optimal equilibria to be arbitrarily large in both FS and CS. Thus, depending on which equilibria are realized, both approaches can be arbitrarily better for either side of the market.
@@
A formal proof can be found in online Appendix B.4. It proceeds by constructing an example with two child and family types where child $c_{1}$ is significantly preferred over $c_{2}$ by all families, while all children slightly prefer $f_{1}$ over $f_{2}$. This implies that family $f_{2}$ is only matched with child $c_{1}$ if no $f_{1}$ family is currently present. The higher search costs in FS then can make $f_{2}$ lose interest in $c_{1}$, regardless of patience levels, causing $f_{2}$ to settle for $c_{2}$. This allows $c_{2}$ to be matched. Conversely, in CS, family $f_{2}$ does not incur high search costs for waiting until they are matched with a $c_{1}$. If they are patient enough, $f_{2}$ therefore prefers waiting for their preferred choice $c_{1}$. This leaves $c_{2}$ without any family interested in them and, therefore, unmatched.
@@
A formal proof can be found in online Appendix B.5. Just as children can benefit from families that decide to settle for a less preferred child, so can other families. It can be the case that a family $f$ is interested in a child $c$ in a CSE but $f$ is not interested in $c$ in an FSE, because the associated expected costs would be too high. Not having $f$ as competition might be enough incentive for another family $f^{\prime}$ to be interested in $c$ under the FSE. As a result, $f^{\prime}$ can be strictly better off in FS.
@@ 6 Effects of Model Parameters
We showed that FSEs cannot be Pareto improvements over CSEs, but CSEs can be Pareto improvements over FSEs. Additionally, we found that some agents can be better off under an FSE compared to a CSE. In order to better understand the conditions under which one of the two approaches might be preferable, we explore the effects that different parameters have on equilibrium outcomes in FS and CS. We provide two more results in favor of CS: First, we show that as families’ patience decreases, at some point all children will be weakly better off in any CSE compared to any FSE. Second, increasing supply on the family side, i.e., increasing the market thickness indicator $\lambda$, can negatively affect children’s utilities in FS but not in CS. Finally, as a sanity check, we investigate the effect of certain parameters or parameter combinations in the limit; all of these latter results can be found in Section D.
@@ 6.1 Discount Factors
For this subsection, let $S^{CS}(\delta_{F}^{\prime})$ and $S^{FS}(\delta_{F}^{\prime})$ denote the set of CSEs and FSEs when $\delta_{F}=\delta_{F}^{\prime}$, respectively. We now show that as families’ patience decreases below a certain threshold, all children will always be better off in CS compared to FS.
@@
A formal proof can be found in online Appendix B.6. Intuitively, the statement follows because in CS, families being interested in a very unlikely match incurs them lower search costs than in FS. While patient families may still not be interested in some children in CS that they are interested in under FS (which drives Proposition 5.7), any sufficiently impatient family will be unwilling to wait.
However, as can be seen in the proof of Proposition 5.8, an analogous statement for families’ utilities and children’s patience level does not hold. Intuitively, a family $f$ might be worse off in CS, because another family $f^{\prime}$ is not shying away from $c$, as $f^{\prime}$ does not have to worry about accumulating wasted search efforts in CS.
@@ 6.2 Market Thickness
Adoption agencies might intuitively prefer to have a larger pool of available families to choose from. Here, we present another result which suggests that this might be generally good in CS but not always in FS when it comes to children’s utilities. Increasing supply on the family side, i.e., increasing the market thickness indicator $\lambda$, can negatively affect children’s utilities in FS but not in CS. For the remainder of Section 6.2, assume that all instance parameters are fixed except for $\lambda$. Let $s^{co-CS,\lambda}$ denote the child-optimal CSE given market thickness indicator $\lambda$. Definitions for $s^{fo-CS,\lambda}$, $s^{co-FS,\lambda}$, and $s^{fo-FS,\lambda}$ are analogous. Proposition 6.2 shows that increasing $\lambda$ can lead to some children being worse off in FS.
@@
A formal proof can be found in online Appendix B.7. Effectively, what is happening is that if multiple families are interested in a child, then increased market thickness $\lambda$ increases competition for the child and therefore the search costs for the less preferred families. If the child is close to indifferent between families, but some families lose interest due to the higher cost, then the resulting decrease in the child’s utility can be larger than the increase caused by a higher chance to match with a slightly more preferred family.
In CS, on the other hand, increasing $\lambda$ can only have a positive effect on children’s utilities in equilibrium outcomes.
@@
A formal proof can be found in online Appendix B.11. The reason why this holds in CS is that, unlike in FS, families will not shy away from children in whom they are interested just because the probability of matching with them decreases. This result is reminiscent of a similar result in standard two-sided matching markets, as the number of agents in one side increases, the other side agents become all unambiguously better off under side-optimal stable matchings (Gale and Sotomayor 1985). However, it only holds for caseworker-driven search and only for the children’s welfare.
@@ 7 Numerical Evaluation
We previously established that CSEs can be Pareto improvements over FSEs while FSEs cannot be Pareto improvements over CSEs, and that agents can be better off in either approach (see Theorem 5.5 and Proposition 5.6). Additionally, we have shown that all children will be better off in CS compared to FS if families are sufficiently impatient. Here, we present numerical results to further investigate under which conditions children and families will be better off in CS or FS. Our results suggest that CS is almost always preferable for both sides of the market. Only when agents’ preferences are perfectly correlated and families are very patient, we find that on average there are more children types better off in FS compared to CS in equilibrium.
@@ 7.1 Setup
Valuations: For the generation of agents’ valuations, we follow other approaches from the matching literature (Abdulkadiroğlu et al. 2015, Mennle et al. 2015). Each agent type $i\in A$ is uniformly assigned a “quality” $q_{i}$ at random from $[0,1]$. Then, for each child-family pair $(c,f)\in C\times F$, idiosyncratic values $\eta_{c}(f)$ and $\eta_{f}(c)$ are randomly drawn from $[0,1]$. For a given value $\alpha\in[0,1]$, we obtain the preliminary valuations $v^{\prime}_{c}(f)=\alpha q_{f}+(1-\alpha)\eta_{c}(f)$ and $v^{\prime}_{f}(c)=\alpha q_{c}+(1-\alpha)\eta_{f}(c)$. Note that as $\alpha$ increases, agents’ preferences become more similar and end up being identical (vertical) for $\alpha=1$. Final valuations $v$ are obtained by normalizing $v^{\prime}$, such that the minimal and maximal value that each agent has for a match is 00 and $1$, respectively.
Data: We generated $200$ quality-value pairs $(q^{(1)},\eta^{(1)}),\ldots,(q^{(200)},\eta^{(200)})$ as described above. Parameters $p$ and $\lambda$ are chosen to be $p=\lambda=0.5$, and we let $\delta:=\delta_{C}=\delta_{F}$ and $\kappa:=\kappa_{C}=\kappa_{F}$ For each pair $(q^{(k)},\eta^{(k)})$, we computed the child-optimal CSE/FSE and the family-optimal CSE/FSE for each combination of $\alpha$, $\delta$, and $\kappa$, where $\alpha\in\{0,1/3,2/3,1.0\}$, $\delta\in\{0.8,0.9,0.975,0.99\}$, and $\kappa\in\{0.01,0.02,0.05,0.1\}$. Thus, we consider $200\cdot 4\cdot 4\cdot 4=12800$ different instances and compute a total of $51200$ equilibria.
@@ 7.2 Equilibrium Computation
The mapping $T$ defined in the proof of Proposition 4.5 can be used to find the child-optimal and family-optimal equilibria. The following procedure converges to an equilibrium threshold profile: Start from the $\leq_{C}$-minimal element in $Y$ and recursively apply $T$ to it. This produces a sequence $y^{0},y^{1},y^{2},\ldots$ of threshold profiles, which converges to the fo-FSE. Starting from the $\leq_{C}$-maximal element yields the co-FSE. In order to terminate after a finite number of steps, we force the procedure to stop once $|y^{k}_{i}-y^{k+1}_{i}|\leq\epsilon$ for all $i\in A$ for some previously chosen small parameter $\epsilon>0$. The threshold profiles obtained by this procedure can then be mapped to the corresponding strategy profiles. In our numerical experiments, we have performed additional checks to confirm that the computed strategy profiles are in fact equilibria.
@@ 7.3 Results
Before comparing FS and CS, we first note that family and child optimal equilibria in FS coincide roughly $97\%$ of the time. The same holds for CS. For simplicity, we only consider family-optimal equilibria in our analysis. As equilibria are almost always unique within each search technology, results for child-optimal equilibria do not differ markedly even with substantial differences between FS and CS. Of all cases considered, the CSE and the FSE only coincide once in the sense that the same agents are mutually interested in each other.
@@
The previous Pareto comparison only allows for a very high-level comparison of FS and CS. In order to better understand the conditions under which certain agents benefit from FS or CS, we evaluate how many agents are better off in CS and how many agents are better off in FS. Our numerical experiments show that all families are almost always better off in CS. We refer the reader to Section E.1 of the e-companion for more details on families’ statistics. For children, the combination of model parameters affects which approach appears more appealing. Figure 2 shows how many children are (strictly) better off (in terms of utilities) in CS and FS for different parameter combinations.
Figure 2: The ratio of children who are on average (strictly) better off in terms of utilities in either approach in the family-optimal equilibrium for different combinations of agents’ patience and preference correlation level.
CS provides higher utility than FS for almost all children when agents are sufficiently impatient (e.g., $\delta=0.8$) because CS allows agents to express interest in more potential match partners without risking wasted search efforts. Being interested in more agents increases the probability of getting matched at each time step, which is especially valuable to children when patience is low. On the other hand, FS incentivizes agents to focus on a smaller set of match candidates due to higher expected total search costs. When $\delta=0.8$, families will on average be interested in $35.9$ and $43.5$ child types in FS and CS, respectively. [^7] The probability of matches occurring is another metric of interest to stakeholders. Our findings on match probabilities can be found in Section E.2. Interestingly, more children benefit from FS than CS when agents are extremely patient and agents’ preferences are almost completely aligned. This explains why CSEs are less frequently Pareto improvements over FSEs under these conditions, as we previously saw in Figure 1.
Figure 3: The ratio of children who are on average (strictly) better off in terms of (expected discounted) match value in either approach in the family-optimal equilibrium for different combinations of agents’ patience and the level of preference correlation.
Figure 3 shows that CS not only reduces wasted search efforts in many cases but also enables children to match with more preferred families. We calculate a child’s match value as the child’s utility ignoring the expected search costs, which might be less relevant to a policymaker trying to improve child outcomes.
@@ 8 Empirical Evidence from a Field Implementation
To validate our model and understand the real-world implications of switching from an FS to a CS approach, we analyze children’s outcomes for a multi-county region in Florida that implemented a CS approach on July 1, 2018, by adopting a technology platform developed by a nonprofit organization. The agency’s previous search approach relied on regular email announcements to a pool of registered families to advertise children in need of adoptive families. In the new approach, the platform allows prospective adoptive parents to complete a questionnaire. Its singular role is to recommend families for a child’s caseworker to consider. The region, one of 20 “circuits” in Florida, is administered by a non-profit agency under contract with the state. Out of frustration with the difficulties in finding adoptive placements for children, the leadership of the non-profit child welfare agency for this circuit decided to implement the technology platform and follow a caseworker-driven search strategy.
This analysis compares the performance of the region’s outcomes for children from the Adoption and Foster Care Analysis and Recruiting System (AFCARS) from FY2015-FY2021, which is the most recently available data on both foster care cases and finalized adoptions. To assess performance, we aggregate individual-level predictions and outcomes for the agency’s children listed on the platform. While the agency listed all or nearly all children on the platform — which is unique among agencies implementing that platform — a small number of children also found placement through other channels, such as word of mouth within the agency, Florida’s online photolisting websites, and other contracted recruitment efforts for specific children. Including these placements helps to understand whether the population of waiting children is better served once an agency implements the platform.
@@
Figure 4: Children eligible for adoption in Florida on July 1 of each year and whether they are adopted by June 30 of the following year (Florida Department of Children and Families 2019, 2023).
@@
Florida experienced at least two dramatic shocks to its child welfare system in the years over which the platform was implemented. As shown in Figure 4, the statewide population of children legally free for adoption — which also includes children on a path to adoption by relatives and foster parents — increased by nearly 50% from 2015 to 2018, possibly due to social challenges such as the opioid epidemic (Florida Department of Children and Families 2019, 2023). Despite this increase in children needing adoptive placements, the state’s reported metric of the percentage of children eligible for adoption on June 30 of some year and adopted by July 1 of the following year remained above 55% from 2017 until 2019. However, that metric dropped below 50% in the years since the coronavirus pandemic, and the last reported value of 46.7% for 2022-2023 is the lowest of all years considered. Thus, maintaining pre-pandemic outcome levels may constitute success for the region implementing the platform.
8.1 Data
Our analysis relies on multiple data sets: the AFCARS Foster Care 6-month File (Children’s Bureau, Administration on Children, Youth and Families 2023b), the AFCARS Adoption File (Children’s Bureau, Administration on Children, Youth and Families 2023a), and case history data from the platform. Regrettably, and despite significant efforts, the agency implementing the platform could not extract information from Florida’s statewide case management system and assemble its own comparable case history dataset, although we provide validation in Section 8.3 that uses circuit-level data in state reports. Filtering children based on case goals and the relationships with adoptive families proved to be the biggest obstacles to using the agency’s data. However, the agency was helpful in manually tracking down the outcomes of children who left the platform without an adoptive placement. From 766,527 foster care 6-month update records, we identified 10,284 children as legally free and clear for adoption with cases starting after October 1, 2014, which is the closest federal fiscal year cut-off for the data in Figure 4. Children’s timelines begin with the termination of parental rights (TPR) order and end with the last status update, which could be the adoption finalization date. To qualify for this analysis, children had to have a case goal of adoption in their final record or previously had a case goal of adoption with (a) a resulting non-relative adoption, (b) a final case goal of “emancipation”, or (c) a discharge reason of “emancipation”. We excluded children listed in the Adoption File as being adopted by a relative, step-parent, or foster parent. Because the most recent Adoptions File only covers adoptions through September 30, 2021, we excluded case data past that date as we would not be able to tell if adoptions resulted from relative, step-parent, or foster placements. We also excluded any cases with a duration less than 120 days — a 30-day appeal window after TPR before a search can start plus a legal minimum of 90 days between placement with the adoptive family and adoption finalization — or over 18 years.
We note that the AFCARS datasets’ inability to explicitly identify children in need of adoptive resources results in a conservative benchmark on the platform’s performance; i.e., it will make the platform appear less helpful than it actually was. Specifically, some children included in the AFCARS dataset may quickly find placements with non-relatives, such as teachers, church members, or neighbors. Such children have short times until adoption but are typically never listed on the platform. While we filter out cases with implausibly short durations of 120 days or fewer, some of these cases with durations exceeding 120 days may still be included in the AFCARS benchmark model, as we cannot identify them separately. Child welfare professionals have also mentioned the possibility that users of the AFCARS reporting system may inadvertantly classify adoptions by foster parents as non-relative adoptions. These potential limitations show the importance of more explicitly identifying children for whom an active search is being conducted in child welfare data sources to aid in evaluating performance.
We received case data from the platform about 369 children in need of adoptive placements who were listed by the agency. The platform provided its first matches around July 1, 2018. Of these children listed on the platform, 176 had finalized adoptions by February 1, 2023. Of the finalized adoptions, the adoptive parents for 143 children were found through the platform. We excluded an additional 66 children who were listed at one time on the platform but achieved permanency through adoptive placement with relatives or foster care parents or were reunified. Our data only includes activity and case updates through February 1, 2023; children adopted after that date are not counted as adopted in our dataset. For example, on February 1, 2023, 25 children were in placements intended to lead to an adoption, but that had not yet been finalized.
8.2 Results
Table 1: Summary Statistics of AFCARS and Platform Data
AFCARS (N=10,284) Platform (N=279)

### v4

@@ 6.2 No Approach Dominates the Other
A formal proof can be found in online Appendix D.4. It proceeds by constructing examples where child and family utilities are mismatched, causing the utility gap between child-optimal and family-optimal equilibria to be arbitrarily large in both FS and CS. Thus, depending on which equilibria are realized, both approaches can be arbitrarily better for either side of the market.
@@
A formal proof can be found in online Appendix D.5. It proceeds by constructing an example with two child and family types where child $c_{1}$ is significantly preferred over $c_{2}$ by all families, while all children slightly prefer $f_{1}$ over $f_{2}$. This implies that family $f_{2}$ is only matched with child $c_{1}$ if no $f_{1}$ family is currently present. The higher search costs in FS can then make $f_{2}$ lose interest in $c_{1}$, regardless of patience levels, causing $f_{2}$ to settle for $c_{2}$. This allows $c_{2}$ to be matched. Conversely, in CS, family $f_{2}$ does not incur high search costs for waiting until they are matched with a $c_{1}$. If they are patient enough, $f_{2}$ therefore prefers waiting for their preferred choice $c_{1}$. This leaves $c_{2}$ without any interested family and, therefore, unmatched.
@@
A formal proof can be found in online Appendix D.6. Just as children can benefit from families that decide to settle for a less preferred child, so can other families. It can be the case that a family $f$ is interested in a child $c$ in a CSE but $f$ is not interested in $c$ in an FSE, because the associated expected costs would be too high. Not having $f$ as competition might be enough incentive for another family $f^{\prime}$ to be interested in $c$ under the FSE. As a result, $f^{\prime}$ can be strictly better off in FS.
@@ 7 Effects of Model Parameters
We showed that FSEs cannot be Pareto improvements over CSEs, but CSEs can be Pareto improvements over FSEs. Additionally, we found that some agents can be better off under an FSE compared to a CSE. In order to better understand the conditions under which one of the two approaches might be preferable, we explore the effects that different parameters have on equilibrium outcomes in FS and CS. We provide two more results in favor of CS: First, we show that as families’ patience decreases, at some point all children will be weakly better off in any CSE compared to any FSE. Second, increasing supply on the family side, i.e., increasing the market thickness indicator $\lambda$, can negatively affect children’s utilities in FS but not in CS. Finally, as a sanity check, we investigate the effect of certain parameters or parameter combinations in the limit; all of these latter results can be found in Section F.
@@ 7.1 Discount Factors
For this subsection, let $S^{CS}(\delta_{F}^{\prime})$ and $S^{FS}(\delta_{F}^{\prime})$ denote the set of CSEs and FSEs when $\delta_{F}=\delta_{F}^{\prime}$, respectively. We now demonstrate that as families’ patience decreases below a certain threshold, all children will always be better off in CS than in FS.
@@
A formal proof can be found in online Appendix D.7. Intuitively, the statement follows because in CS, families’ interest in a very unlikely match incurs lower search costs than in FS. While patient families may still not be interested in some children in CS that they are interested in under FS (which drives Proposition 6.6), any sufficiently impatient family will be unwilling to wait.
However, as can be seen in the proof of Proposition 6.7, an analogous statement for families’ utilities and children’s patience level does not hold. Intuitively, a family $f$ might be worse off in CS, because another family $f^{\prime}$ is not shying away from $c$, as $f^{\prime}$ does not have to worry about wasted search efforts in CS.
@@ 7.2 Market Thickness
Adoption agencies might intuitively prefer to have a larger pool of available families to choose from. Here, we present another result that suggests this might be generally beneficial in CS but not always in FS when it comes to children’s utilities. Increasing supply on the family side, i.e., increasing the market thickness indicator $\lambda$, can negatively affect children’s utilities in FS but not in CS. For the remainder of Section 7.2, assume that all instance parameters are fixed except for $\lambda$. Let $s^{co-CS,\lambda}$ denote the child-optimal CSE given market thickness indicator $\lambda$. Definitions for $s^{fo-CS,\lambda}$, $s^{co-FS,\lambda}$, and $s^{fo-FS,\lambda}$ are analogous. Proposition 7.2 shows that increasing $\lambda$ can lead to some children being worse off in FS.
@@
A formal proof can be found in online Appendix D.8. Effectively, what is happening is that if multiple families are interested in a child, then increased market thickness $\lambda$ increases competition and, therefore, the search costs for less preferred families. If the child is close to being indifferent between families and some families lose interest due to the higher cost, then the resulting decrease in the child’s utility can be larger than the increase caused by a higher chance to match with a slightly more preferred family.
In CS, on the other hand, increasing $\lambda$ can only have a positive effect on children’s utilities in equilibrium.
@@
A formal proof can be found in online Appendix D.12. The reason why this holds in CS is that, unlike in FS, families will not shy away from children in whom they are interested just because the probability of matching with them decreases. This result is reminiscent of a similar result in standard two-sided matching markets, as the number of agents in one side increases, the other side agents become all unambiguously better off under side-optimal stable matchings (Gale and Sotomayor 1985). However, it only holds for CS and only for the children’s welfare.
@@ 8 Numerical Evaluation
We previously established that CSEs can be Pareto improvements over FSEs while FSEs cannot be Pareto improvements over CSEs, and that agents can be better off in either approach (see Theorem 6.4 and Proposition 6.5). Additionally, we have shown that all children will be better off in CS compared to FS if families are sufficiently impatient. Here, we present numerical results to further investigate the conditions under which children and families will be better off in CS or FS. Our results suggest that CS is almost always preferable for both sides of the market. Only when agents’ preferences are perfectly correlated and families are very patient do we find that, on average, there are more child types better off in FS than in CS in equilibrium.
@@ 8.1 Setup
Valuations: For the generation of agents’ valuations, we follow other approaches from the matching literature (Abdulkadiroğlu et al. 2015, Mennle et al. 2015). Each agent type $i\in A$ is uniformly assigned a “quality” $q_{i}$ at random from $[0,1]$. Then, for each child-family pair $(c,f)\in C\times F$, idiosyncratic values $\eta_{c}(f)$ and $\eta_{f}(c)$ are randomly drawn from $[0,1]$. For a given value $\alpha\in[0,1]$, we obtain the preliminary valuations $v^{\prime}_{c}(f)=\alpha q_{f}+(1-\alpha)\eta_{c}(f)$ and $v^{\prime}_{f}(c)=\alpha q_{c}+(1-\alpha)\eta_{f}(c)$. Note that as $\alpha$ increases, agents’ preferences become more similar and end up being identical (vertical) for $\alpha=1$. Final valuations $v$ are obtained by normalizing $v^{\prime}$, such that the minimal and maximal value that each agent has for a match is 00 and $1$, respectively. Although we consider various values for $\alpha$, note that the practical level of verticality in child welfare tends to be relatively low. For example, data provided by the platform that we examine in Section 9 shows notable variation in families’ stated preferences on several dimensions, which, taken together, suggests that preferences are not very aligned. See online Appendix G for details.
Data: We generated $200$ quality-value pairs $(q^{(1)},\eta^{(1)}),\ldots,(q^{(200)},\eta^{(200)})$ as described above. Parameters $p$ and $\lambda$ are chosen to be $p=\lambda=0.5$, and we let $\delta:=\delta_{C}=\delta_{F}$ and $\kappa:=\kappa_{C}=\kappa_{F}$ For each pair $(q^{(k)},\eta^{(k)})$, we computed the child-optimal CSE/FSE and the family-optimal CSE/FSE for each combination of $\alpha$, $\delta$, and $\kappa$, where $\alpha\in\{0,1/3,2/3,1.0\}$, $\delta\in\{0.8,0.9,0.975,0.99\}$, and $\kappa\in\{0.01,0.02,0.05,0.1\}$. Thus, we consider $200\cdot 4\cdot 4\cdot 4=12800$ different instances and compute a total of $51200$ equilibria. Unless specified otherwise, results are averaged over all instances.
@@ 8.2 Equilibrium Computation
The mapping $T$ defined in the proof of Proposition 5.5 can be used to find the child-optimal and family-optimal equilibria. The following procedure converges to an equilibrium threshold profile: Start from the $\leq_{C}$-minimal element in $Y$, i.e., the minimum point of the lattice spanned by the partial order $\leq_{C}$ over the threshold vectors in the game, and recursively apply $T$ to it. The $\leq_{C}$-minimal element in $Y$ is the threshold vector $\underline{y}$ where $\underline{y}_{c}=0$ for each child $c$ and $\underline{y}_{f}=\bar{v}$ for each family $f$. This produces a sequence $y^{0},y^{1},y^{2},\ldots$ of threshold profiles, which converges to the fo-FSE. Starting from the $\leq_{C}$-maximal element in $Y$, i.e., the maximum point of the lattice spanned by the partial order $\leq_{C}$ over the threshold vectors in the game, yields the co-FSE. The $\leq_{C}$-maximal element is the threshold vector $\overline{y}$ where $\overline{y}_{c}=\bar{v}$ for each child $c$ and $\overline{y}_{f}=0$ for each family $f$. In order to terminate after a finite number of steps, we force the procedure to stop once $|y^{k}_{i}-y^{k+1}_{i}|\leq\epsilon$ for all $i\in A$ for some previously chosen small parameter $\epsilon>0$. The threshold profiles obtained by this procedure can then be mapped to the corresponding strategy profiles. We have performed additional checks to validate that the computed strategy profiles are indeed equilibria.
@@ 8.3 Results
Before comparing FS and CS, we first note that family- and child-optimal equilibria in FS coincide roughly $97\%$ of the time. The same holds for CS. For simplicity, we only consider family-optimal equilibria in our analysis. As equilibria are almost always unique within each search technology, results for child-optimal equilibria do not differ markedly. We do, however, observe substantial differences between FS and CS. Of all cases considered, the CSE and the FSE only coincide once in the sense that the same agents are mutually interested in each other.
@@
The previous Pareto comparison only allows for a very high-level comparison of FS and CS. To better understand the conditions under which certain agents benefit from FS or CS, we compare the number of agents who are better off in either search discipline. Our numerical experiments show that all families are almost always better off in CS. We refer the reader to online Appendix H.1 for more details on families’ statistics. For children, the combination of model parameters affects which approach appears more appealing. Figure 2 shows how many children are (strictly) better off (in terms of utilities) in CS and FS for different parameter combinations.
Figure 2: The ratio of children on average (strictly) better off in terms of utilities in either approach in the family-optimal equilibrium for different combinations of agents’ patience and preference correlation level.
CS provides higher utility than FS for almost all children when agents are sufficiently impatient (e.g., $\delta=0.8$) because CS allows agents to express interest in more potential match partners without risking wasted search efforts. Being interested in more agents increases the probability of getting matched at each time step, which is especially valuable to children when patience is low. On the other hand, FS incentivizes agents to focus on a smaller set of match candidates due to higher expected total search costs. When $\delta=0.8$, families will, on average, be interested in $35.9$ and $43.5$ child types in FS and CS, respectively. [^10] The probability of matches occurring is another metric of interest to stakeholders. Our findings on match probabilities can be found in Section H.2. Interestingly, more children benefit from FS than CS when agents are extremely patient and agents’ preferences are almost completely aligned. Although unlikely to occur in practical child welfare settings, this explains why CSEs are less frequently Pareto improvements over FSEs under these conditions, as we previously saw in Figure 1.
Figure 3: The ratio of children on average (strictly) better off in terms of (expected discounted) match value in either approach in the family-optimal equilibrium for different combinations of agents’ patience and the level of preference correlation.
Figure 3 shows that CS not only reduces wasted search efforts in many cases but also enables children to match with more preferred families. We calculate a child’s match value as the child’s utility, ignoring the expected search costs, which might be less relevant to a policymaker trying to improve child outcomes.
@@ 9 Empirical Evidence from a Field Implementation
To validate our model and understand the real-world implications of switching from an FS to a CS approach, we analyze children’s outcomes for a multi-county region in Florida that started implementing a CS approach on July 1, 2018, by adopting a technology platform which we simply refer to as the platform. The region, one of 20 circuits in Florida, is administered on the state’s behalf by a non-profit organization — which we refer to as the agency. The agency’s previous FS approach relied on regular email announcements to the full pool of registered families to advertise children in need of adoption. Out of frustration with the difficulties in finding adoptive placements for children, the agency’s leadership decided to implement the technology platform and follow a CS discipline. In the new approach, caseworkers contact individual families listed on the platform whom they consider a good match for a child. The platform’s purpose in this approach is to provide an easy-to-navigate database of prospective families. Additionally, it offers family recruitment, evaluation, and recommendation services to support caseworkers.
@@
Figure 4: Children eligible for adoption in Florida on July 1 of each year and whether they are adopted by June 30 of the following year (Florida Department of Children and Families 2019, 2024).
Our analysis compares the agency’s performance to outcomes for children in Florida reported in the Adoption and Foster Care Analysis and Recruiting System (AFCARS) from the US Children’s Bureau, Administration on Children, Youth and Families (2023a, b) for FY2015 to FY2021, which is the most recently available data on all foster care cases and finalized adoptions. From 766,527 AFCARS foster care 6-month update records for Florida children during this period, we identified 10,286 children as legally free and clear for adoption with cases starting after October 1, 2014. While the platform’s usage is free to all agencies in Florida, the overwhelming majority of placements during the observed period were still searched via traditional methods, with only the considered agency utilizing CS as their primary method. While some other agencies occasionally employed the platform’s tools to help identify families for harder-to-place children, these constitute less than $4\%$ of adoptions in the AFCARS dataset.
For the agency, we obtained case data from the platform about 279 children in need of adoptive placements who were listed by the agency before the AFCARS cut-off date of October 1, 2021. The platform provided its first matches around July 1, 2018. While the agency listed all children in need of an adoptive placement on the platform, a small number of these children found placement through other channels, such as word of mouth within the agency, Florida’s online photolisting websites, and other contracted recruitment efforts for specific children. We include these placements in our analysis because serendipitous matches occur outside the primary (FS or CS) search channel in any system but cannot be identified in the AFCARS dataset. Details of how the data is assembled and pre-processed can be found in online Appendix I.1.
@@
It should be noted that Florida experienced at least two dramatic shocks to its child welfare system in the years over which the platform was implemented. As shown in Figure 4, the statewide population of children legally free for adoption — which also includes children on a path to adoption by relatives and foster parents — increased by nearly 50% from 2015 to 2018 (Florida Department of Children and Families 2019, 2024). Quast et al. (2018) document a relationship between opioid prescriptions and child welfare system entries in Florida in the early 2010s that could partially explain this increase in children in need of adoption as the opioid crisis worsened. Despite an increase in children needing adoptive placements, the state’s reported metric of the percentage of children eligible for adoption on June 30 of some year and adopted by July 1 of the following year remained above 55% from 2017 until 2019. However, that metric dropped below 50% in the years after the coronavirus pandemic, as caseworker turnover, staffing shortages, and slower judicial processing times may have hindered adoptive searches. After reaching the lowest reported value of 46.7% for 2022-2023, the metric only returned to above 50% in the most recent report covering 2023-2024.
9.1 Statistical Approaches
Using the AFCARS and platform datasets, we assess the platform’s impact on children’s outcomes through two statistical approaches that take advantage of the datasets’ similar structure. First, we construct a benchmark for the set of agency children using the Florida AFCARS case data. Second, we statistically measure the treatment effect of the platform by appending the platform case data to the AFCARS case data and including a variable for being listed on the platform. This approach dampens the estimated platform effect and its statistical significance, as the 279 children of the platform are double-counted in the AFCARS dataset without the treatment variable. Thus, we expect real treatment effects to be larger than those estimated.
Table 1 presents child statistics for the AFCARS and platform data, highlighting some key differences: The platform’s population of children tends to be older, more male, and more likely to have a clinically diagnosed disability — all factors associated with greater difficulty in placing children. The difference in the age distribution between the AFCARS and platform datasets is statistically significant at the $0.05$ level by both a Welch’s t-test and a Mann–Whitney U test. Using a chi-squared test, the difference is also significant at that level for the Female variable but not for the Disability variable. Fewer children on the platform are Black, which reflects regional variations in Florida’s population. Due to differences in how the datasets treat multi-racial children, we provide analysis in Section I.2.3 of the Appendix that shows how estimates of the platform’s performance improve if the Black variable refers only to children with the Black or African American variable exclusively selected as a race variable in the AFCARS dataset.
We also note that the AFCARS dataset’s inability to explicitly identify children in need of adoptive search resources results in a conservative assessment of the platform’s performance; i.e., it will underestimate the platform’s impact. Specifically, managers with experience in Florida child welfare agencies report that some children in the AFCARS dataset may go through the termination of parental rights (TPR) judicial process with an already-identified adoptive placement with a non-relative, such as a teacher, church member, or neighbor.
Table 1: Summary Statistics for AFCARS and Platform Data
AFCARS (N=10,286) Platform (N=279)